If y(x-1)=z then x=
I REALLY NEED SOME HELP WITH THIS ONE, PLEASE THANKS.

Answer by checkley71(8403)

Y(X-1)=Z
XY-Y=Z
XY=Y+Z
X=(Y+Z)/Y ANSWER.